OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 2x2 + 9y2 = 18

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

\[\frac{x^2}{9}+\frac{y^2}{2}=1\\ a=3\qquad b=\sqrt{2} \]

OpenStudy (anonymous):

rest follows

OpenStudy (anonymous):

Is this the answer then? Center: (0, 0); Vertices: (0, -3), (0, 3); Foci: (0, -\[\sqrt{7}\]

OpenStudy (anonymous):

and (0, \[\sqrt{7}\] )

OpenStudy (anonymous):

\[e=\sqrt{1-(2/9)}=\sqrt{7}/3\]

OpenStudy (anonymous):

\[f=ea=\sqrt{7}\] yep all well

OpenStudy (anonymous):

Ok thanks so much for your help!